if a population has a standard deviation of 20, what is the standard error of the mean if the sample size 16 is selected? 


1
Expert's answer
2021-05-23T22:12:05-0400

"SEM=\frac{\sigma}{\sqrt{n}}=\frac{20}{\sqrt{16}}=5."